The mean number of students ill at a school is 3.8 per day, for the first 20 school days of a term. On the 21st day 8 students are ill. What is the mean after 21 days? 


1
Expert's answer
2020-11-04T15:38:08-0500

Total for the first 20 days

"\bar x = 3.8" and "n=20"


"n \bar x = 20*3.8 =76"

Total for 21 days


"\frac{76+8}{21} = 4"

The mean after 21 days is 4